# Paying and Persuading

### v4

@@ 7 Discussion
$V_{+}^{\prime}(\mathcal{U}(\{h^{\zeta},\mu,a\}|\sigma^{*}))=-k$, the logic of the pullback lemma implies there is a $\mathcal{U}(\{h^{\zeta},\mu,a\}|\sigma^{*})$-constrained optimum, $\tilde{\sigma}$, which gives full information and pays Receiver an additional amount (over the potential payoff from the pullback) to meet the utility promise constraint which is also optimal. Modifying $\sigma^{*}$ so that $\sigma^{*}=\bar{\sigma}$ at any history which is not succeeded by $h^{\zeta}$ on the path of play, and otherwise replacing the continuation strategy $\sigma^{*}(\cdot|h^{\zeta})$ with $\bar{\sigma}(\cdot|h^{\zeta})=\tilde{\sigma}$ delivers the desired continuation strategy $\bar{\sigma}$ with $h^{\zeta}\not\in\mathcal{H}^{\zeta}(\bar{\sigma})$.
Second, suppose $\eta=u^{RFI}(\mu)-\mathcal{U}(\{h^{\zeta},\mu,a\}|\sigma^{*})>0$. Take $\tilde{u}=\frac{\delta}{1-\delta}t(h^{\zeta};\sigma)(\mu,a)$. If $\eta>\tilde{u}$, then replace $t(h^{\zeta})(\mu,a)$ with 0, but replace $\mathcal{U}(\{h^{\zeta},\mu,a\})$ with $\tilde{u}$ exactly. Note Sender is indifferent to this change since $V_{+}^{\prime}(\mathcal{U}(\{h^{\zeta},\mu,a))=-k$. The replacement procedure in the first case and the fourth property of the pullback lemma then imply there exists a sequentially optimal continuation strategy starting from $h^{\zeta}$ at which payments are 0 under $(\mu,a)$. If instead $\eta<\tilde{u}$, set $t(h^{\zeta};\bar{\sigma})(\mu,a)=t(h^{\zeta};\sigma)(\mu,a)-\frac{\delta}{1-\delta}\eta$ and set continuation strategies to the obedient $\sigma^{*}$ in the pullback lemma. Because $\tilde{u}>\eta$, set $\alpha=1$ in the fourth property of the pullback lemma, i.e. there is full revelation at all future continuation histories. As before, this maintains optimality of the continuation strategy without affecting the joint distribution of actions in histories before $s$ (or incomparable to $h^{\zeta}$). In either case, transfers are now a last resort at $h^{\xi}$ conditional on $(\mu,a)$. Repeating this procedure for all $(\mu,a)$ that appear on path implies the argument. ∎
@@
$\mathbb{E}_{\mu}[v(a^{\prime},\theta)-v(a,\theta)]\geq\min_{\theta\in\Theta}v(a^{\prime},\theta)-v(a,\theta)\geq k\max_{\theta\in\Theta}u(a,\theta)-u(a^{\prime},\theta)\geq k\mathbb{E}_{\mu}[u(a,\theta)-u(a^{\prime},\theta)].$
@@
From here, fix some $\beta\in(0,1)$ be chosen so that $\beta\delta^{s-t}\eta\mathbb{P}^{\sigma}(h^{\zeta}|(h^{\xi},\mu))\leq t(h^{\xi};\sigma)(\mu)$. Consider the alternative strategy $\tilde{\sigma}$, which modifies $\sigma$ in the following way:
@@
    At $h^{\xi}$, $t(h^{\xi};\tilde{\sigma})(\mu,a)=t(h^{\xi};\sigma)(\mu,a)-\beta\delta^{s-t}\eta\mathbb{P}^{\sigma}(h^{\zeta}|(h^{\xi},\mu))$. At any continuation strategy $h^{\phi}\succ h^{\xi}$ where $h^{\phi}\not\succsim h^{\zeta}$, do not change the strategy from $\sigma$.
@@
    At $h^{\zeta}$, with probability $\beta$, set $\sigma^{S}(h^{\zeta})=\{\tilde{\rho},\tilde{t}\}$ with continuation payoffs $\tilde{u}^{\prime}$, and set $\sigma^{R}(h^{\zeta})=\tilde{\alpha}(h^{\zeta})$. With complementary probability $1-\beta$, do not change the continuation strategy; that is, $\sigma^{S}(\cdot|h^{\zeta})=\tilde{\sigma}^{S}(\cdot|h^{\zeta})$.
@@
This observation then implies Sender’s payoff changes only at two histories: $h^{\xi}$ and $h^{\zeta}$. At $h^{\xi}$, they pay $k\beta\delta^{s-t}\eta\mathbb{P}^{\sigma}(h^{\zeta}|(h^{\xi},\mu))$ less in transfers, while at $h^{\zeta}$ they forego $\beta\varepsilon$ total payoff in order to do so. This changes their total expected payoff by
$\delta^{\xi}\mathbb{P}^{\tilde{\sigma}}(h^{\xi})\left(k\beta\delta^{s-t}\eta\mathbb{P}^{\sigma}(h^{\zeta}|(h^{\xi},\mu))+\beta\delta^{s-t}\varepsilon\mathbb{P}^{\tilde{\sigma}}(h^{\zeta}|(h^{\xi},\mu))\right).$
@@
Fix some $(\gamma,m)\in\Gamma^{int}(\mu^{\infty})$. Let $\sigma^{*}$ be the stationary strategy where $\rho$ reveals the state and recommends mixed action $\gamma_{\theta}=\mathbb{E}[\gamma|\theta]\in\Delta(A)$ at $\delta_{\theta}$ at every history on-path, and pays $m$. If Receiver deviates from any recommendation, then give no information forevermore and recommend Receiver’s no-information optimum. At any history, any state $\theta$, and any action $a$, Receiver’s deviation gain is bounded from above by
@@
$(1-\delta)(u(a,\theta)+m)+\delta\left(\mathbb{E}_{\gamma}[u(a,\theta)]+m\right)\geq(1-\delta)u(a^{*}(\delta_{\theta}),\theta)+\delta\sum_{\xi=1}^{\infty}u(a^{*}(M^{\xi}\delta_{\theta},\mathbf{0}),\theta).$
@@
Consider the experiment $\tilde{\rho}_{i}$ which splits beliefs along $\tilde{\mu}_{i}$ in the following way. First, for any $\mu_{i}\neq\tilde{\mu}_{i}$, let $\rho_{i}(\mu_{i})=\tilde{\rho}_{i}(\mu_{i})$. Along $\tilde{\mu}_{i}$, split beliefs so that $\tilde{\rho}_{i}(\tilde{\mu}_{i})=0$, and $\tilde{\rho}_{i}(\delta_{0})=\rho_{i}(\tilde{\mu}_{i}(0))$, and $\tilde{\rho}_{i}(\delta_{1})=\rho_{i}(\tilde{\mu}_{i}(1))$. This is clearly Bayes plausible. In a similar way, modify payments and utility promises so that they are the same at beliefs where $\rho_{i}=\tilde{\rho}_{i}$, and equal to $t_{i}(\tilde{\mu}_{i})$ and $u_{i}^{\prime}(\tilde{\mu}_{i})$ at the newly induced beliefs. Finally, allow Receiver to take her favorite action at the newly induced beliefs, and otherwise the same action as before.
Such an experiment must be incentive compatible because Receiver is myopically maximizing at the new beliefs, and at the old ones (i.e. those supported by $\rho$), they are the same. Do not change any $\rho_{j}$, $j\neq i$. This alternative strategy strictly increases $\mathbb{P}(a_{1}=1|\theta_{1}=1,h^{\xi},\sigma)$ while decreasing $\mathbb{P}(a_{1}=0|\theta_{1}=1,h^{\xi},\sigma)$ without affecting decisions in any other dimension. This increases Sender’s stage-game payoff. Moreover, it does not change Sender’s expected payments or continuation value (as the distribution of $\tilde{t}$ and $\tilde{u}^{\prime}$ are the same as $t$ and $u^{\prime}$). Thus this deviation is strictly profitable, contradicting supposed optimality of $\sigma$. ∎
@@
We have $V_{i+}^{\prime}(u)=-c_{i}\text{ if }u\leq u^{FI},c_{i}<k$, and $V_{i+}^{\prime}(u)=-k$ otherwise.
@@
There exists an optimal strategy $\sigma$ where for $\mathbb{P}^{\sigma}$ almost all histories $h^{\xi}$,
@@
$\mathcal{G}_{i}(h,\sigma)=\#\{t:\rho(h^{\xi};\sigma)\not\in\{\rho^{BP},\rho^{FI}\}\text{ and }\mathcal{G}_{i}(\sigma)=\{h:\mathcal{G}_{i}(h,\sigma)\geq 2\}.$
@@
Now fix some on-path history $h\in\mathcal{G}_{i}(\sigma)\cap\mathbb{P}^{\sigma}$, so that for $t=\tilde{T}_{i}^{*}(h;\sigma)$, one has $\rho_{i}(h^{\phi};\sigma)\not\in\{\rho^{BP},\rho^{FI}\}$ for $r=t,s$ where $s\geq t$, $h^{\zeta}\succ h^{\xi}$. Suppose $\text{supp}(\rho(h^{\xi};\sigma))=\{0,\mu_{i}(h^{\xi})\}$ and $\text{supp}(\rho(h^{\zeta};\sigma))=\{0,\mu_{i}(h^{\zeta})\}$. Consider now small $\varepsilon<\mu_{i}(h^{\xi})-\frac{1}{2}$ and experiments $\tilde{\rho}_{i}(h^{\xi};\sigma)$ and $\tilde{\rho}_{i}(h^{\zeta};\sigma)$ such that $\text{supp}(\tilde{\rho}_{i}(h^{\xi};\sigma)))=\{0,\mu_{i}(h^{\xi})-\varepsilon\}$ and $\text{supp}(\tilde{\rho}_{i}(h^{\zeta};\sigma))=\{0,\mu_{i}(h^{\zeta})+\eta(\varepsilon)\}$ for some $\eta(\varepsilon)$ vanishing as $\varepsilon\to 0$. By Lemma 7, both of these changes are experiments that remain on the Pareto frontier. By Lemma 8, there is moreover some $\eta(\varepsilon)$ such that by changing $\rho_{i}(h^{\phi};\sigma)\to\tilde{\rho}_{i}(h^{\phi};\sigma)$ for $\phi\in\{\zeta,\xi\}$ and identifying the utility promises after $\mu_{i}$ to the newly induced beliefs (i.e. changing nothing else), one remains on the Pareto frontier and in particular promises the same amount of utility to the agent at all histories.
@@
Appendix B: The Strong Topology on Strategies
In this appendix we formally define the strong topology for strategies $\sigma:H^{\infty}\to\mathcal{X}^{\infty}$ where $\mathcal{X}$ is a normed compact topological space, $H$ is some finite alphabet [^19] I remark here that the space of histories in the paper is not drawn from some finite alphabet since $\Delta(\Theta)\times A$ is uncountable. However, by a revelation principle, the space of strategies is isomorphic (up to payoffs) to one where $M=A$ (though taking this isomorphism immediately complicates the proof of Theorem 1 and hence is not invoked explicitly). However, when $M=A$, the alphabet can be taken to be exactly $A$, which is finite. Thus the strong topology can be invoked in the proof of Theorem 1. and $H^{\infty}$ and $\mathcal{X}^{\infty}$ denote the product topological spaces over $H$ and $\mathcal{X}$, respectively. It will be useful to identify strategies $\sigma$ with their basis decompositions, i.e. write it as a sequence of outputs of time-$\xi$ histories where the $t$-th coordinate is the element of $\mathcal{X}$ mapped by the time-$\xi$ projection of histories, i.e.
@@
To give an economic interpretation, say $H^{\infty}$ is the set of infinite histories, $h^{\xi}$ is the set of time-$\xi$ histories, $H$ is the alphabet of public signals, and $\mathcal{X}$ is the (potentially mixed) set. $\sigma$ is then a strategy in the corresponding repeated game. This is the context in which these objects are used in this paper, and a version of this topology appears in Luo and Wolitzky (2025)’s Lemma 3. In this appendix, we formally prove and collect several relevant definitions and facts about this topological space. Throughout, endow $H$ with the discrete metric and metrize $\mathcal{X}$ with its norm denoted by $||\cdot||_{\infty}$.
@@
and so in particular if any two histories $h,h^{\prime}$ agree for the first $\frac{1}{2^{T}}$ periods (in which case the distance $d(h,h^{\prime})$ is at most $\frac{1}{2^{T-1}}$) then it must be that the distance of their images is separated by at most $\frac{1}{2^{T-1}}$. Since this is true for any $\sigma$ and at any history, the modulus of continuity for functions in $\Sigma$ is independent both of the choice of $h$ and of the choice of $\sigma$ and hence $\Sigma$ is equicontinuous. Moreover, this space of strategies is closed; for any sequence $\{\sigma_{n}\}$ which converges to some $\sigma$, $\sigma$ is itself a function defined on the cylinders as
@@
where convergence of $\sigma_{n}(h^{\xi})$ is guaranteed as $\mathcal{X}$ is compact. Finally, for any fixed $h\in H$, the closure of the set $\{\sigma(h)\}_{\sigma\in\Sigma}$ is a closed subset of a compact set (namely, $\mathcal{X}^{\infty}$) and hence compact. This implies we satisfy Theorem 47.1 of Munkres (2000) and hence $\Sigma$ is sequentially compact (though it is proven in the topology of compact convergence—this coincides with our topology when the ambient space of the domain is compact). ∎
Now suppose $\mathcal{X}=\Delta(H)$, so that each time-$\xi$ history is mapped to a (random) new letter of the alphabet. Then for any any function $\sigma$, note that there is a natural measure defined on $\mathcal{H}^{\infty}$ generated by the cylinders $\mathbb{P}_{\xi}^{\sigma}(h^{\xi})=\mathbb{P}_{\xi-1}^{\sigma}(h^{\xi-1})\sigma(h^{\xi-1})(x)$, where $h^{\xi}=\{h^{\xi-1},x\}$ for $x\in H$. Let the tail probability measure associated with this sequence be $\mathbb{P}^{\sigma}$ (that is, $\mathbb{P}^{\sigma}(h)=\lim\limits_{\xi\to\infty}\mathbb{P}_{\xi}^{\sigma}(h)$). We conclude with the following result.
@@
If $\sigma_{n}\to\sigma$ in the strong topology, $\mathbb{P}^{\sigma_{n}}\to\mathbb{P}^{\sigma}$ pointwise.

### v5

@@ 7 Discussion
$V_{+}^{\prime}(\mathcal{U}(\{h^{\zeta},\mu,a\}|\sigma^{*}))=-k$, the logic of the pullback lemma implies there is a $\mathcal{U}(\{h^{\zeta},\mu,a\}|\sigma^{*})$-constrained optimum, $\tilde{\sigma}$, which gives full information and pays Receiver an additional amount (over the potential payoff from the pullback) to meet the utility promise constraint which is also optimal. Modifying $\sigma^{*}$ so that $\sigma^{*}=\bar{\sigma}$ at any history which is not succeeded by $h^{\zeta}$ on the path of play, and otherwise replacing the continuation contract $\sigma^{*}(\cdot|h^{\zeta})$ with $\bar{\sigma}(\cdot|h^{\zeta})=\tilde{\sigma}$ delivers the desired continuation contract $\bar{\sigma}$ with $h^{\zeta}\not\in\mathcal{H}^{\zeta}(\bar{\sigma})$.
Second, suppose $\eta=u^{RFI}(\mu)-\mathcal{U}(\{h^{\zeta},\mu,a\}|\sigma^{*})>0$. Take $\tilde{u}=\frac{1-\delta}{\delta}t(h^{\zeta};\sigma)(\mu,a)$. If $\eta>\tilde{u}$, then replace $t(h^{\zeta})(\mu,a)$ with 0, but add to $\mathcal{U}(\{h^{\zeta},\mu,a\})$ the amount $\tilde{u}$ exactly. Note Sender is indifferent to this change since $V_{+}e^{\prime}(\mathcal{U}(\{h^{\zeta},\mu,a))=-k$. The replacement procedure in the first case and the fourth property of the pullback lemma then imply there exists a sequentially optimal continuation contract starting from $h^{\zeta}$ at which payments are 0 under $(\mu,a)$. If instead $\eta<\tilde{u}$, set $t(h^{\zeta};\bar{\sigma})(\mu,a)=t(h^{\zeta};\sigma)(\mu,a)-\frac{\delta}{1-\delta}\eta$ and set continuation contracts to be the obedient $\sigma^{*}$ in the pullback lemma. Because $\tilde{u}>\eta$, set $\alpha=1$ in the fourth property of the pullback lemma, i.e. there is full revelation at all future continuation histories. As before, this maintains optimality of the continuation contract without affecting the joint distribution of actions in histories before $s$ (or incomparable to $h^{\zeta}$). In either case, transfers are now a last resort at $h^{\xi}$ conditional on $(\mu,a)$. Repeating this procedure for all $(\mu,a)$ that appear on path implies the argument. ∎
@@
$\mathbb{E}_{\mu}[v(a^{\prime},\theta)-v(a,\theta)]\geq\min_{\theta\in\Theta}v(a^{\prime},\theta)-v(a,\theta)>k\max_{\theta\in\Theta}u(a,\theta)-u(a^{\prime},\theta)\geq k\mathbb{E}_{\mu}[u(a,\theta)-u(a^{\prime},\theta)].$
@@
From here, fix some $\beta\in(0,1)$ be chosen so that $\beta\delta^{\zeta-\xi}\eta\mathbb{P}^{\sigma}(h^{\zeta}|(h^{\xi},\mu))\leq t(h^{\xi};\sigma)(\mu)$. Consider the alternative contract $\tilde{\sigma}$, which modifies $\sigma$ in the following way:
@@
    At $h^{\xi}$, $t(h^{\xi};\tilde{\sigma})(\mu,a)=t(h^{\xi};\sigma)(\mu,a)-\beta\delta^{\zeta-\xi}\eta\mathbb{P}^{\sigma}(h^{\zeta}|(h^{\xi},\mu))$. At any continuation contract $h^{\phi}\succ h^{\xi}$ where $h^{\phi}\not\succsim h^{\zeta}$, do not change the contract from $\sigma$.
@@
    At $h^{\zeta}$, with probability $\beta$, set $\sigma^{S}(h^{\zeta})=\{\tilde{\rho},\tilde{t}\}$ with continuation payoffs $\tilde{u}^{\prime}$, and set $\sigma^{R}(h^{\zeta})=\tilde{\alpha}(h^{\zeta})$. With complementary probability $1-\beta$, do not change the continuation contract; that is, $\sigma^{S}(\cdot|h^{\zeta})=\tilde{\sigma}^{S}(\cdot|h^{\zeta})$.
@@
This observation then implies Sender’s payoff changes only at two histories: $h^{\xi}$ and $h^{\zeta}$. At $h^{\xi}$, they pay $k\beta\delta^{\zeta-\xi}\eta\mathbb{P}^{\sigma}(h^{\zeta}|(h^{\xi},\mu))$ less in transfers, while at $h^{\zeta}$ they forego $\beta\varepsilon$ total payoff in order to do so. This changes their total expected payoff by
$\delta^{\xi}\mathbb{P}^{\tilde{\sigma}}(h^{\xi})\left(k\beta\delta^{\zeta-\xi}\eta\mathbb{P}^{\sigma}(h^{\zeta}|(h^{\xi},\mu))+\beta\delta^{\zeta-\xi}\varepsilon\mathbb{P}^{\tilde{\sigma}}(h^{\zeta}|(h^{\xi},\mu))\right).$
@@
Define first the value function $V^{0}$ to be the value function obtained from (FE) with the additional restriction that transfers must always be 0, i.e. there are no transfers. This is the pure dynamic information design problem. Note the domain of $V(\cdot,\mu_{0},\delta)$ is $0,u^{RFI}(\mu_{0})]$. Define $\bar{k}=\inf_{\mu_{0}\in\Delta(\Theta)}V_{-}^{0^{\prime}}(u^{RFI}(\mu_{0}))$ to be a lower bound on the left derivative of $V^{0}$ at its right endpoint $u^{RFI}(\mu_{0})$. Suppose now the result is false, so there exists $k>\bar{k}$ and optimal $\sigma$ at primitives $\{u,v,\delta,k\}$ such that transfers are not a last resort. This implies there exists $h^{\zeta}\succ(h^{\xi},\mu,a)$ on the path of play such that $\rho(h^{\zeta};\sigma)\neq\rho^{FI}$ where $t(h^{\xi};\sigma)(\mu,a)>0$.
Let $\mu^{\zeta}$ be the belief Receiver has history $h^{\zeta}$. By mechanically pushing transfers back into history $h^{\zeta}$, the fact $t(h^{\xi};\sigma)(\mu,a)>0$ implies there exists solution $\{\rho,t,u^{\prime},\alpha\}\in\mathcal{F}(\mathcal{U}(h^{\zeta};\sigma),\mu^{\zeta})$ where $\rho\not\in\rho^{FI}$ and $t\neq 0$ everywhere. Yet because transfers were used at $h^{\xi}$ preceding $h^{\zeta}$, $V_{+}^{\prime}(\mathcal{U}(h^{\zeta};\sigma))=-k$ by Lemma 3. Yet this contradicts the fact $\rho\neq\rho^{FI}$, since $V_{+}^{0}(\cdot,\mu^{\zeta})\geq-\bar{k}>-k$, so just providing informational incentives must be more efficient than k-k. In particular, this implies decreasing transfers $t$ towards 0, and increasing $\rho$ towards $\rho^{FI}$ must be a profitable improvement over $\{\rho,t,u^{\prime},\alpha\}$, contradicting the assumption that tuple was optimal. ∎
PROOF OF PROPOSITION 9
Proof.
$t^{I}(\mu,a)=\mathbb{E}_{\mu}[u(a^{*}(\mu,\mathbf{0}),\theta)-u(a,\theta)]\mathbf{1}\left\{a\in a^{S}(\mu)\right\}$
where
$a^{S}(\mu)=\operatorname*{\arg\!\max}_{a\in A}\mathbb{E}_{\mu}[v(a,\theta)-k(u(a^{*}(\mu,\mathbf{0}),\theta)-u(a,\theta))].$
Lemma 7.
Proof.
$\mathbb{E}_{\mu}[u(a^{*}(\mu,t^{*}),\theta)]+t^{*}(\mu,a^{*}(\mu,t^{*}),)\geq\mathbb{E}_{\mu}[u(a,\theta)]+t^{*}(\mu,a)\text{ for all }a\in A$
Taking differences implies
$\mathbb{E}_{\mu}[u(a^{*}(\mu,t^{*}),\theta)-u(a,\theta)]\geq t^{*}(\mu,a^{*}(\mu,t^{*}))-t^{*}(\mu,a)\text{ for all }a\in A$
$\mathbb{E}_{\mu}[u(a^{*}(\mu,t^{*}),\theta)-u(a,\theta)]\geq\mathbb{E}_{\mu}[u(a^{*}(\mu,t^{*}),\theta)-u(a^{*}(\mu,\mathbf{0})),\theta)]=t^{I}(\mu,a^{*}(\mu,t^{I}))$
Given that canonical transfers are sufficient, it is clear that
$\max_{t}\mathbb{E}_{\mu}[v(a^{*}(\mu,t),\theta)-kt(\mu,a))]=\max_{a\in A}\mathbb{E}_{\mu}[v(a,\theta)+k(u(a,\theta)-u(a^{*}(\mu,\mathbf{0}),\theta))]$
To do so, we need to prove a quick geometric lemma.
Lemma 8.
$\mathcal{O}_{a}$ is a convex polytope.
Proof.
$V^{t}(\mu)=\max_{a\in A}\left\{\mathbb{E}_{\mu}[v(a,\theta)-k(u(\tilde{a},\theta)-u(a,\theta))]\right\},$
PROOF OF PROPOSITION 10
Proof.
@@
Fix some $(\gamma,m)\in\Gamma^{int}(\mu^{\infty})$. Let $\sigma^{*}$ be the stationary contract where $\rho$ reveals the state and recommends mixed action $\gamma_{\theta}=\mathbb{E}[\gamma|\theta]\in\Delta(A)$ at $\delta_{\theta}$ at every history on-path, and pays $m$. If Receiver deviates from any recommendation, then give no information forevermore and recommend Receiver’s no-information optimum. At any history, any state $\theta$, and any action $a$, Receiver’s deviation gain is bounded from above by
@@
$(1-\delta)(u(a,\theta)+m)+\delta\left(\mathbb{E}_{\gamma}[u(a,\theta)]+m\right)\geq(1-\delta)u(a^{*}(\delta_{\theta},\mathbf{0}),\theta)+\sum_{\xi=1}^{\infty}\delta^{\xi}u(a^{*}(M^{\xi}\delta_{\theta},\mathbf{0}),\theta).$
@@
Consider the experiment $\tilde{\rho}_{i}$ which splits beliefs along $\tilde{\mu}_{i}$ in the following way. First, for any $\mu_{i}\neq\tilde{\mu}_{i}$, let $\rho_{i}(\mu_{i})=\tilde{\rho}_{i}(\mu_{i})$. Along $\tilde{\mu}_{i}$, split beliefs so that $\tilde{\rho}_{i}(\tilde{\mu}_{i})=0$, and $\tilde{\rho}_{i}(\delta_{0})=\rho_{i}(\tilde{\mu}_{i})\tilde{\mu}_{i}(0)$, and $\tilde{\rho}_{i}(\delta_{1})=\rho_{i}(\tilde{\mu}_{i})\tilde{\mu}_{i}(1)$. This is clearly Bayes plausible. In a similar way, modify payments and utility promises so that they are the same at beliefs where $\rho_{i}=\tilde{\rho}_{i}$, and equal to $t_{i}(\tilde{\mu}_{i})$ and $u_{i}^{\prime}(\tilde{\mu}_{i})$ at the newly induced beliefs. Finally, allow Receiver to take her favorite action at the newly induced beliefs, and otherwise the same action as before.
Such an experiment must be incentive compatible because Receiver is myopically maximizing at the new beliefs, and at the old ones (i.e. those supported by $\rho$), they are the same. Do not change any $\rho_{j}$, $j\neq i$. This alternative contract strictly increases $\mathbb{P}(a_{1}=1|\theta_{1}=1,h^{\xi},\sigma)$ while decreasing $\mathbb{P}(a_{1}=0|\theta_{1}=1,h^{\xi},\sigma)$ without affecting decisions in any other dimension. This increases Sender’s stage-game payoff. Moreover, it does not change Sender’s expected payments or continuation value (as the distribution of $\tilde{t}$ and $\tilde{u}^{\prime}$ are the same as $t$ and $u^{\prime}$). Thus this deviation is strictly profitable, contradicting supposed optimality of $\sigma$. ∎
@@
We have $V_{i+}^{\prime}(u)=-c_{i}\text{ if }u\leq u^{RFI}(\mu_{0}),c_{i}<k$, and $V_{i+}^{\prime}(u)=-k$ otherwise.
@@
There exists an optimal contract $\sigma$ where for $\mathbb{P}^{\sigma}$ almost all histories $h^{\xi}$,
@@
$\mathcal{G}_{i}(h,\sigma)=\#\{t:\rho(h^{\xi};\sigma)\not\in\{\rho^{BP},\rho^{FI}\}\}\text{ and }\mathcal{G}_{i}(\sigma)=\{h:\mathcal{G}_{i}(h,\sigma)\geq 2\}.$
@@
Now fix some on-path history $h\in\mathcal{G}_{i}(\sigma)\cap\mathbb{P}^{\sigma}$, so that for $\xi=\tilde{T}_{i}^{*}(h;\sigma)$, one has $\rho_{i}(h^{\phi};\sigma)\not\in\{\rho^{BP},\rho^{FI}\}$ for $\phi=\xi,\zeta$ where $\zeta\geq\xi$, $h^{\zeta}\succ h^{\xi}$. Suppose $\text{supp}(\rho(h^{\xi};\sigma))=\{0,\mu_{i}(h^{\xi})\}$ and $\text{supp}(\rho(h^{\zeta};\sigma))=\{0,\mu_{i}(h^{\zeta})\}$. Consider now small $\varepsilon<\mu_{i}(h^{\xi})-\frac{1}{2}$ and experiments $\tilde{\rho}_{i}(h^{\xi};\sigma)$ and $\tilde{\rho}_{i}(h^{\zeta};\sigma)$ such that $\text{supp}(\tilde{\rho}_{i}(h^{\xi};\sigma)))=\{0,\mu_{i}(h^{\xi})-\varepsilon\}$ and $\text{supp}(\tilde{\rho}_{i}(h^{\zeta};\sigma))=\{0,\mu_{i}(h^{\zeta})+\eta(\varepsilon)\}$ for some $\eta(\varepsilon)$ vanishing as $\varepsilon\to 0$. By Lemma 9, both of these changes are experiments that remain on the Pareto frontier. By Lemma 10, there is moreover some $\eta(\varepsilon)$ such that by changing $\rho_{i}(h^{\phi};\sigma)\to\tilde{\rho}_{i}(h^{\phi};\sigma)$ for $\phi\in\{\zeta,\xi\}$ and identifying the utility promises after $\mu_{i}$ to the newly induced beliefs (i.e. changing nothing else), one remains on the Pareto frontier and in particular promises the same amount of utility to the agent at all histories.
@@
Appendix B: The Strong Topology on Contracts
In this appendix I formally define the strong topology for contracts $\sigma:H^{\infty}\to\mathcal{X}^{\infty}$ where $\mathcal{X}$ is a compact subset of a normed topological space [^22] Note by the Kuratowski embedding theorem assuming the existence of the norm is without loss of generality, because every metric space can be isometrically embedded in a normed space., $H$ is some finite alphabet [^23] I remark here that the space of histories in the paper is not drawn from some finite alphabet since $\Delta(\Theta)\times A$ is uncountable. To apply the strong topology, apply the following transformation. First, reduce $\Sigma$ to equivalence classes modulo $\mathbb{Q}^{\sigma}$, $\Sigma/\mathbb{Q}^{\sigma}$: that is, $\sigma\equiv\sigma^{\prime}$ if $\mathbb{Q}^{\sigma}=\mathbb{Q}^{\sigma^{\prime}}$. Second, use the revelation principle to write the space of contracts $\tilde{\Sigma}$ where $S=A$ directly; this is without loss of generality up to $\mathbb{Q}^{\sigma}$. Thus $\tilde{\Sigma}/\mathbb{Q}^{\sigma}=\Sigma/\mathbb{Q}^{\sigma}$, and we can endow the latter with the strong topology as we can endow the former with the strong topology. Thus the strong topology can be invoked in the proof of Theorem 1 on the space of equivalence classes up to $\Sigma/\mathbb{Q}^{\sigma}$, which is sufficient for our purposes to apply the squeezing lemma and conclude that transfers are a last resort. and $H^{\infty}$ and $\mathcal{X}^{\infty}$ denote the product topological spaces over $H$ and $\mathcal{X}$, respectively. It will be useful to identify contracts $\sigma$ with their basis decompositions, i.e. write it as a sequence of outputs of time-$\xi$ histories where the $t$-th coordinate is the element of $\mathcal{X}$ mapped by the time-$\xi$ projection of histories, i.e.
@@
To give an economic interpretation, say $H^{\infty}$ is the set of infinite histories, $h^{\xi}$ is the set of time-$\xi$ histories, $H$ is the alphabet of public signals, and $\mathcal{X}$ is the (potentially mixed) action set. $\sigma$ is then a strategy in the corresponding repeated game. This is the context in which these objects are used in this paper, and a version of this topology appears in Luo and Wolitzky (2025)’s Lemma 3. In this appendix, I formally prove and collect several relevant definitions and facts about this topological space. Throughout, endow $H$ with the discrete metric and metrize $\mathcal{X}$ with its norm denoted by $||\cdot||_{\infty}$.
@@
and so in particular if any two histories $h,h^{\prime}$ agree for the first $T$ periods (in which case the distance $d(h,h^{\prime})$ is at most $\frac{1}{2^{T-1}}$) then it must be that the distance of their images is separated by at most $\frac{1}{2^{T-1}}$. Since this is true for any $\sigma$ and at any history, the modulus of continuity for functions in $\Sigma$ is independent both of the choice of $h$ and of the choice of $\sigma$ and hence $\Sigma$ is equicontinuous. Moreover, this space of contracts is closed; for any sequence $\{\sigma_{n}\}$ which converges to some $\sigma$, $\sigma$ is itself a function defined on the cylinders as
@@
where convergence of $\sigma_{n}(h^{\xi})$ is guaranteed as $\mathcal{X}$ is compact. Finally, for any fixed $h\in H^{\infty}$, the closure of the set $\{\sigma(h)\}_{\sigma\in\Sigma}$ is a closed subset of a compact set (namely, $\mathcal{X}^{\infty}$) and hence compact. Note also $H^{\infty}$ is compact as well. This implies we satisfy Theorem 47.1 of Munkres (2000) and hence $\Sigma$ is sequentially compact (though it is proven in the topology of compact convergence—this coincides with our topology when the ambient space of the domain is compact). ∎
Now suppose $\mathcal{X}=\Delta(H)$, so that each time-$\xi$ history is mapped to a (random) new letter of the alphabet. Then for any any function $\sigma$, note that there is a natural measure defined on $\mathcal{H}^{\infty}$ generated by the cylinders $\mathbb{P}_{\xi}^{\sigma}(h^{\xi})=\mathbb{P}_{\xi-1}^{\sigma}(h^{\xi-1})\sigma(h^{\xi-1})(x)$, where $h^{\xi}=\{h^{\xi-1},x\}$ for $x\in H$. Let the Kolmogorov extension of the consistent family $\{P_{\xi}^{\sigma}\}$ to the cylinder $\sigma$-algebra be $\mathbb{P}^{\sigma}$. I conclude with the following result.
@@
If $\sigma_{n}\to\sigma$ in the strong topology, $\mathbb{P}^{\sigma_{n}}\to\mathbb{P}^{\sigma}$ pointwise (i.e. on cylinder sets).